OpenStudy (anonymous):

For f(x)=log base 3 of (x-2), find f-1(x).

OpenStudy (anonymous):

\[f(x)=\log_3(x-2)\] \[y=\log_3(x-2)\] now switch x and y and solve for y \[x=\log_3(y-2)\] equivalent exponential form is \[3^x=y-2\] add 2 and get \[y=3^x+2\] or \[f^{-1}(x)=3^x+2\]

OpenStudy (anonymous):

great. thanks